OpenStudy (anonymous):

find the mean, median, and mode of the two data sets. 1. 5 7 10 10 7 9 5 6 7 7 4 2. 12 13 14 9 7 8 15 10 11 6

OpenStudy (anonymous):

@hoblos mind helping me with this? :D

OpenStudy (hoblos):

i would like to.. but actually i forgot the formulas of these!!

OpenStudy (anonymous):

@hoblos http://www.purplemath.com/modules/meanmode.htm

OpenStudy (hoblos):

for the first one: first rearrange 4 5 5 6 7 7 7 7 9 10 10 we have 11numbers mean = (4+5+5+6+7+7+7+7+9+10+10)/11 = 77/11 = 7 median = 6th number =7 mode = 7 (repeated 4 times)